# Bellman Optimality of Average-Reward Robust Markov Decision Processes with a Constant Gain

### v1

@@
[…]
Bellman Optimality of Average-Reward Robust Markov Decision Processes with a Constant Gain
   Shengbo Wang
Daniel J. Epstein Department of
Industrial and Systems Engineering
University of Southern California
shengbow@usc.edu
&Nian Si
Department of Industrial Engineering
and Decision Analytics
Hong Kong University of Science and Technology
niansi@ust.hk
@@ 1 Introduction
In this paper, we consider optimality conditions for average-reward robust Markov decision processes (MDPs), where a controller aims to optimize the long-run average reward in the presence of an adversary. Specifically, we consider the value
@@
$\overline{\alpha}(\mu,\Pi,\mathrm{K}):=\sup_{\pi\in\Pi}\inf_{\kappa\in\mathrm{K}}\limsup_{n\rightarrow\infty}E_{\mu}^{\pi,\kappa}\frac{1}{n}\sum_{k=0}^{n-1}r(X_{k},A_{k}),$
@@
where $\Pi$ and $\mathrm{K}$ denote the controller’s and adversary’s policy classes, respectively, and $\mu$ is the initial distribution. We assume that the adversary is S-rectangular, and that both policy classes may be history-dependent (denoted by $\Pi_{\mathrm{H}}$ and $\mathrm{K}_{\mathrm{H}}$) or stationary (denoted by $\Pi_{\mathrm{S}}$ and $\mathrm{K}_{\mathrm{S}}$), with potentially asymmetric information structures between the controller and adversary decisions.
@@
We are interested in identifying conditions under which $\overline{\alpha}(\mu,\Pi,\mathrm{K})=\alpha^{*}$ holds for all initial distributions $\mu$. Here, $\alpha^{*}\in[0,1]$ arises as part of a solution pair $(u^{*},\alpha^{*})$ to the following robust Bellman equation with a constant gain:
$u^{*}(s)=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}E_{\phi,p_{s}}\!\left[r(s,A_{0})-\alpha^{*}+u^{*}(X_{1})\right],$ (1.1)
where $\mathcal{Q}$ and $\mathcal{P}_{s}$ are the controller’s and adversary’s decision sets, respectively.
Our main results can be summarized as follows.
@@
  • Theorem 4 in Section 4 implies that if the controller is communicating (as in Definition 3) and $\mathcal{Q}$ is compact, then (1.1) has a solution.
  • Theorem 5 in Section 4 shows that if the adversary is communicating, $\mathcal{P}_{s},\;s\in S$ are convex and compact, and $\mathcal{Q}$ is convex, then the Bellman equation (1.1) has a solution.
@@ 1.1 Literature review
Robust MDPs: While the Bellman optimality of discounted-reward robust MDPs has been extensively studied [5, 9, 10, 7], the corresponding results for the average-reward setting remain underexplored. To the best of our knowledge, Wang et al., 2023b [8] provides the first results under the stringent assumptions of SA-rectangularity and uniform unichains. Grand-Clement et al., [4] focus on Blackwell optimality, showing that $\epsilon$-Blackwell optimal policies always exist under SA-rectangularity. Moreover, they demonstrate that in S-rectangular RMDPs, average-optimal policies may fail to exist; and even when they do exist, they may need to be strictly history-dependent.
@@
Stochastic games (SGs): S-rectangular robust MDPs can be viewed as a generalization of zero-sum SGs. They extend the standard SGs framework in the following ways: (i) an asymmetry of information, where the controller may use history-dependent policies while the adversary is restricted to stationary or Markovian ones, and (ii) ambiguity-set constraints, where the adversary’s feasible set may be nonconvex, in contrast to the convex mixed-strategy sets of SGs.
There is an extensive literature on average-payoff SGs; see Section 5 of Filar and Vrieze, [3]. The present work, which establishes constant-gain Bellman optimality for robust MDPs under one-sided communication and compactness assumptions, also contributes to the SG literature.
@@ 2 Canonical Construction and the Optimal Robust Control Problem
The history set $\mathbf{H}_{t}$ at time $t$ contains all $t$-truncated sample paths $\mathbf{H}_{t}:=\left\{{h_{t}=(s_{0},a_{0},\dots,a_{t-1},s_{t}):\omega=(s_{0},a_{0},s_{1}\dots)\in\Omega}\right\}.$ We also define the random element $H_{t}:\Omega\rightarrow\mathbf{H}_{t}$ by point evaluation $H_{t}(\omega)=h_{t}$, and the $\sigma$-field $\mathcal{H}_{t}:=\sigma(H_{t})$.
Given a prescribed subset $\mathcal{Q}\subset\mathcal{P}(\mathcal{A})$, a controller policy $\pi$ is a sequence of decision rules $\pi=(\pi_{0},\pi_{1},\pi_{2},\ldots)$ where each $\pi_{t}$ is a measure-valued function $\pi_{t}:\mathbf{H}_{t}\rightarrow\mathcal{Q}$, represented in conditional distribution form as $\pi_{t}(a|h_{t})\in[0,1]$ with $\sum_{a\in A}\pi_{t}(a|h_{t})=1$. The history-dependent controller policy class is therefore $\Pi_{\mathrm{H}}(\mathcal{Q}):=\{\pi=(\pi_{0},\pi_{1},\ldots):\pi_{t}\in\{\mathbf{H}_{t}\rightarrow\mathcal{Q}\},\ \forall t\geq 0\}$.
A controller policy $\pi=(\pi_{0},\pi_{1},\ldots)$ is stationary if for any $t_{1},t_{2}\geq 0$ and $h_{t_{1}}\in\mathbf{H}_{t_{1}},h^{\prime}_{t_{2}}\in\mathbf{H}_{t_{2}}$ such that $s_{t_{1}}=s^{\prime}_{t_{2}}$, we have $\pi_{t_{1}}(\cdot|h_{t_{1}})=\pi_{t_{2}}(\cdot|h^{\prime}_{t_{2}})$. In particular, this means $\pi_{t}(a|h_{t})=\Delta(a|s_{t})$ where $h_{t}=(s_{0},a_{0},\dots,s_{t})$ for some $\Delta:S\to\mathcal{Q}$ for all $t\geq 0$. Thus, a stationary controller policy can be identified with $\Delta:S\to\mathcal{Q}$, i.e., $\pi=(\Delta,\Delta,\ldots)$. Accordingly, the stationary policy class for the controller is $\Pi_{\mathrm{S}}(\mathcal{Q}):=\{(\Delta,\Delta,\ldots):\Delta\in\{S\to\mathcal{Q}\}\}$, which is identified with $\{S\to\mathcal{Q}\}$.
On the adversary side, for each $s\in S$ we fix a prescribed set of measure-valued functions $\mathcal{P}_{s}\subset\{A\to\mathcal{P}(\mathcal{S})\}$. The product set $\mathcal{P}:=\bigtimes_{s\in S}\mathcal{P}_{s}$ is called an S-rectangular ambiguity set.
Given $\{\mathcal{P}_{s}:s\in S\}$, a history-dependent $S$-rectangular adversary policy $\kappa$ is a sequence of adversarial decision rules $\kappa=(\kappa_{0},\kappa_{1},\kappa_{2},\ldots)$. Each decision rule $\kappa_{t}$ specifies the conditional distribution of the next state given a history $h_{t}\in\mathbf{H}_{t}$ and an action $a\in A$, i.e., $\kappa_{t}(s^{\prime}|h_{t},a)\in[0,1]$ with $\sum_{s^{\prime}\in S}\kappa_{t}(s^{\prime}|h_{t},a)=1$. The history-dependent adversary policy class is $\mathrm{K}_{\mathrm{H}}(\mathcal{P}):=\{\kappa=(\kappa_{0},\kappa_{1},\ldots):\kappa_{t}(\cdot|h_{t},\cdot)\in\mathcal{P}_{s_{t}},\ \text{where }h_{t}=(s_{0},a_{0},\ldots,s_{t}),\ \forall t\geq 0\}$.
@@ 3 Robust Bellman Equations and Optimality
In this section, we define the constant-gain robust Bellman equation and show that its solution determines the long-run average reward of the robust control problem. This also implies stationary optimality for the controller in the $\overline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}})$ and $\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}})$ case.
3.1 Robust Bellman Equations with a Constant Gain
@@
The existence and uniqueness of the solution to (3.3) follows from a standard contraction argument (see Wang et al., 2023a [7]). However, existing literature has only shown the existence of a solution to (3.1) in the SA-rectangular settings under further assumptions [8, Theorem 8].

### v2

@@
Abstract
Learning and optimal control under robust Markov decision processes (MDPs) have received increasing attention, yet most existing theory, algorithms, and applications focus on finite-horizon or discounted models. Long-run average-reward formulations, while natural in many operations research and management contexts, remain underexplored. This is primarily because the dynamic programming foundations are technically challenging and only partially understood, with several fundamental questions remaining open. This paper steps toward a general framework for average-reward robust MDPs by analyzing the constant-gain setting. We study the average-reward robust control problem with possible information asymmetries between the controller and an S-rectangular adversary. Our analysis centers on the constant-gain robust Bellman equation, examining both the existence of solutions and their relationship to the optimal average reward. Specifically, we identify when solutions to the robust Bellman equation characterize the optimal average reward and stationary policies, and we provide one-sided weak communication conditions ensuring solutions’ existence. These findings expand the dynamic programming theory for average-reward robust MDPs and lay a foundation for robust dynamic decision making under long-run average criteria in operational environments.
@@ 1 Introduction
Markov Decision Processes (MDPs) provide a foundational framework for modeling sequential decision-making under uncertainty, underpinning much of modern data-driven dynamic decision-making and reinforcement learning (RL) [22]. Data-driven stochastic control continues to advance and attract new research interest [2, 9]. In parallel, the past decade has witnessed remarkable successes of RL algorithms in increasingly sophisticated simulated environments—including superhuman performance in Atari games [16], mastery of Go [20], and progress toward AI reasoning agents [8].
Nevertheless, generalizability and robustness of these methods to out-of-sample, real-world environments remain limited, owing to model misspecification and sim-to-real gaps that can arise from discrepancies in dynamics, partial observability, stochasticity, and unaccounted real-world perturbations. To bridge this gap and enhance policy reliability in practical deployment, the robust MDP framework has emerged as a principled approach, explicitly accounting for model ambiguity and worst-case environment disturbances while still preserving (in most cases) the tractability of MDP models [10, 17, 14, 28, 24].
While discounted-reward robust MDPs have been relatively well-studied, the average-reward setting introduces significant theoretical challenges that remain largely unresolved. In particular, the optimality conditions, commonly referred to as the Bellman optimality equations, have not yet been fully characterized for average-reward robust MDPs. This gap hinders the development of efficient algorithms, sample-efficient learning methods, and performance guarantees, motivating our investigation into these foundational open problems.
@@
In this paper, we consider optimality conditions for average-reward robust Markov decision processes (MDPs). Specifically, we consider the finite-state and finite-action setting and we focus on
@@
$\overline{\alpha}(\mu,\Pi,\mathrm{K}):=\sup_{\pi\in\Pi}\inf_{\kappa\in\mathrm{K}}\limsup_{n\rightarrow\infty}E_{\mu}^{\pi,\kappa}\left[\frac{1}{n}\sum_{k=0}^{n-1}r(X_{k},A_{k})\right],$
@@
where $\Pi$ and $\mathrm{K}$ denote the controller’s and adversary’s policy classes, respectively, and $\mu$ is the initial distribution. We consider S-rectangular adversary policy classes, in which the adversary’s perturbation of the transition probabilities is state-wise separable: the choice at one state does not affect the set of admissible choices at any other state. However, action dependencies within each state may remain coupled, thereby enforcing constraints among the adversary’s choices for different actions at the same state. Moreover, both controller and adversary’s policy classes may be history-dependent (denoted by $\Pi_{\mathrm{H}}$ and $\mathrm{K}_{\mathrm{H}}$) or stationary (denoted by $\Pi_{\mathrm{S}}$ and $\mathrm{K}_{\mathrm{S}}$), with potentially asymmetric information structures between the controller and adversary decisions.
@@
We are interested in the constant-gain Bellman optimality, that is, identifying conditions under which $\overline{\alpha}(\mu,\Pi,\mathrm{K})=\alpha^{*}$ holds for all initial distributions $\mu$. Here, $\alpha^{*}\in[0,1]$ is part of a solution pair $(u^{*},\alpha^{*})$ to the following robust Bellman equation with a constant gain:
$u^{*}(s)=\sup_{\phi\in\mathcal{Q}}\inf_{p_{s}\in\mathcal{P}_{s}}\mathbb{E}_{\phi,p_{s}}\!\left[r(s,A_{0})-\alpha^{*}+u^{*}(X_{1})\right],$ (1.1)
where the expectation is taken w.r.t. the measure $P_{\phi,p_{s}}(A_{0}=a,X_{1}=s^{\prime})=\phi(a)p_{s,a}(s^{\prime})$. Here, $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{A})$ denotes the controller’s admissible decision set at state $s$. In particular, we consider two types of decision sets: (i) the deterministic policy set $\mathcal{Q}=\{\delta_{a}:a\in\mathcal{A}\}$, where $\delta_{a}$ is the Dirac measure at action $a$; and (ii) the fully randomized action set $\mathcal{Q}=\mathcal{A}$. Moreover, $\mathcal{P}_{s}$ denotes the adversary’s decision set, which can be understood as the projection of the ambiguity set $\mathcal{P}$ onto state $s$.
In standard MDP settings, it is well understood that weak communication is sufficient for constant-gain Bellman optimality, and that a solution to the Bellman equation characterizes an optimal stationary deterministic policy [18]. Extending these results to robust MDPs, however, presents significant theoretical challenges for several reasons:
  1. 1.
    The relationship between the solution of the robust Bellman equation and the optimal robust control value is not straightforward. Grand-Clement et al., [7] provides an example, adapted from the classical Big Match game, showing that the optimal control value $\overline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K})$ can be strictly larger than the value obtained from any stationary policy.
  2. 2.
    It is well known that, to achieve optimal decision-making, robust MDPs may require randomized policies [28]. Consequently, there is no reason to expect a randomized policy to be Blackwell optimal [7]. This invalidates a direct application of the standard arguments that establish Bellman optimality in classical MDPs.
  3. 3.
    Weak communication-type assumptions are much harder to deploy and analyze under the controller-adversary dynamics. In particular, different stationary controller/adversary policies can induce different communicating classes. Hence, unlike classical MDPs, it is unclear a priori what the communicating class associated with an optimal policy should be for robust MDPs.
In this paper, we first clarify the implications of a solution to the constant-gain robust Bellman equation (1.1), including the extent to which it characterizes the optimal robust control value and the associated policies. When (1.1) does characterize the optimal robust control, we introduce one-sided (weak) communication conditions–natural generalizations of their classical MDP counterparts to robust MDPs–and show that they are sufficient to ensure the existence of a solution to (1.1). Specifically, the controller is said to be (weakly) communicating if, for every stationary controller policy, the induced MDP faced by the adversary is (weakly) communicating. Likewise, the adversary is (weakly) communicating if every MDP within the ambiguity set $\mathcal{P}$ is (weakly) communicating. For cases in which Bellman optimality fails due to information asymmetry, we provide a thorough treatment that yields a necessary and sufficient characterization of the control value. Our main results can be summarized as follows.
@@
  • Theorem 3, together with Remark 2, certifies the optimality of the policy derived from any solution of the robust Bellman equation.
  • Theorem 5 shows that if the controller is weakly communicating (as in Definition 3) and $\mathcal{Q}$ and $\mathcal{P}_{s}$ are compact for all $s\in S$, then (1.1) has a solution.
  • Theorem 6 and 7 together imply that if the adversary is weakly communicating and $\mathcal{Q}$ and $\mathcal{P}_{s}$ are convex and compact for all $s\in S$, then the Bellman equation (1.1) has a solution.
@@ 1.1 Literature review
Robust MDPs: While the Bellman optimality of discounted-reward robust MDPs has been extensively studied [10, 28, 29, 24], the corresponding results for the average-reward setting remain underexplored. To the best of our knowledge, Wang et al., 2023c [26] provides the first results under the stringent assumptions of SA-rectangularity and uniform unichains. Grand-Clement et al., [7] focus on Blackwell optimality, showing that $\epsilon$-Blackwell optimal policies always exist under SA-rectangularity. Moreover, they demonstrate that in S-rectangular RMDPs, average-reward optimal policies may fail to exist; and even when they do exist, they may need to be strictly history-dependent.
@@
Stochastic games (SGs): S-rectangular robust MDPs can be viewed as a generalization of two-player zero-sum SGs. They extend the standard SGs framework in the following ways: (i) an asymmetry of information, where the controller may use history-dependent policies while the adversary is restricted to stationary or Markovian ones, and (ii) ambiguity-set constraints, where the adversary’s feasible set may be infinite and nonconvex, in contrast to the convex mixed-strategy sets of SGs. Below, we provide a detailed review of what is known in the stochastic game literature.
Mertens and Neyman, [15] establish that $\epsilon$-optimal strategies always exist for all players, implying that every zero-sum stochastic game admits a value in the finite-state and finite-action setting. In the special case of irreducible stochastic games, Section 5 of Filar and Vrieze, [5] shows that both players possess optimal stationary strategies, that the optimal value is state-independent, and that the Bellman equation admits a solution equal to this common value. A concise overview of these foundational results can also be found in the tutorial by Renault, [19]. To further weaken these assumptions, Wei et al., [27] show that when the SG is player 1 communicating, the Bellman equation still admits a solution that characterizes the value of the game. The player 1 communication condition is equivalent to the controller-communication assumption in our paper. However, as in the literature on stochastic games, their players’ policies can be randomized, making the decision sets convex. For continuous-state or continuous-action settings, analogous results have been derived under the geometric ergodicity assumption [13, 11, 12].
Finally, Garrec, [6] study communicating zero-sum product stochastic games, where each player has an individual state that evolves solely according to their own previous state and action. However, their notion of communication differs from the one considered in our formulation. In fact, their setting does not satisfy the (weakly) communicating assumption used in this paper. As shown by Garrec, [6] as well as Vigeral, [23], Sorin and Vigeral, [21], Ziliotto, [31], the value of average-reward stochastic games may fail to exist without compactness or communicating assumptions when the action spaces are infinite.
1.2 Comments on Paper Organization
In the sections that follow, we set the stage and present the paper’s main results. To guide the reader, we provide a brief roadmap and highlight the technical flow.
Section 2 gives a rigorous, self-contained formulation of the controller–adversary dynamics and the optimal control objective in a robust MDP. Section 3 introduces the constant gain robust Bellman equation and, conditional on the existence of a solution, derives its implications for the optimal robust control problem. Section 4 provides sufficient conditions for the existence of solutions to the robust Bellman equation, focusing on one-sided (weak) communication-type structures that arise naturally in classical MDPs. Finally, Section 6 analyzes a special case where information asymmetry forces the optimal robust control value to equal the value of the robust Bellman equation with the $\sup$ and $\inf$ interchanged; in this regime, no stationary policy can be near-optimal unless the original and exchanged equations have the same constant gain.
The main theorems are organized as follows. Theorems 1, 2, and 3 establish consequences assuming that the constant gain robust Bellman equation admits a solution. Theorems 4, 5, 6, and 7, together with their corollaries, give sufficient conditions on the controller’s and adversary’s decision sets that guarantee the existence of a solution. Finally, Theorem 8 treats an asymmetric-information setting in which, even when a solution exists, the optimal robust control value need not coincide with that solution.
@@ 2 Canonical Construction and the Optimal Robust Control Problem
The history set $\mathbf{H}_{t}$ at time $t$ contains all $t$-truncated sample paths
$\mathbf{H}_{t}:=\left\{{h_{t}=(s_{0},a_{0},\dots,a_{t-1},s_{t}):\omega=(s_{0},a_{0},s_{1}\dots)\in\Omega}\right\}.$
We also define the random element $H_{t}:\Omega\rightarrow\mathbf{H}_{t}$ by point evaluation $H_{t}(\omega)=h_{t}$, and the $\sigma$-field $\mathcal{H}_{t}:=\sigma(H_{t})$.
Given a prescribed subset $\mathcal{Q}\subseteq\mathcal{P}(\mathcal{A})$, a controller policy $\pi$ is a sequence of decision rules $\pi=(\pi_{0},\pi_{1},\pi_{2},\ldots)$ where each $\pi_{t}$ is a measure-valued function $\pi_{t}:\mathbf{H}_{t}\rightarrow\mathcal{Q}$, represented in conditional distribution form as $\pi_{t}(a|h_{t})\in[0,1]$ with $\sum_{a\in A}\pi_{t}(a|h_{t})=1$. The history-dependent controller policy class is therefore
$\Pi_{\mathrm{H}}(\mathcal{Q}):=\{\pi=(\pi_{0},\pi_{1},\ldots):\pi_{t}\in\{\mathbf{H}_{t}\rightarrow\mathcal{Q}\},\ \forall t\geq 0\}.$
A controller policy $\pi=(\pi_{0},\pi_{1},\ldots)$ is stationary if for any $t_{1},t_{2}\geq 0$ and $h_{t_{1}}\in\mathbf{H}_{t_{1}},h^{\prime}_{t_{2}}\in\mathbf{H}_{t_{2}}$ such that $s_{t_{1}}=s^{\prime}_{t_{2}}$, we have $\pi_{t_{1}}(\cdot|h_{t_{1}})=\pi_{t_{2}}(\cdot|h^{\prime}_{t_{2}})$. In particular, this means $\pi_{t}(a|h_{t})=\Delta(a|s_{t})$ where $h_{t}=(s_{0},a_{0},\dots,s_{t})$ for some $\Delta:S\to\mathcal{Q}$ for all $t\geq 0$. Thus, a stationary controller policy can be identified with $\Delta:S\to\mathcal{Q}$, i.e., $\pi=(\Delta,\Delta,\ldots)$. Accordingly, the stationary policy class for the controller is
$\Pi_{\mathrm{S}}(\mathcal{Q}):=\{(\Delta,\Delta,\ldots):\Delta\in\{S\to\mathcal{Q}\}\},$
which is identified with $\{S\to\mathcal{Q}\}$.
On the adversary side, for each $s\in S$ we fix a prescribed set of measure-valued functions $\mathcal{P}_{s}\subseteq\{A\to\mathcal{P}(\mathcal{S})\}$. The product set $\mathcal{P}:=\bigtimes_{s\in S}\mathcal{P}_{s}$ is called an S-rectangular ambiguity set.
Given $\{\mathcal{P}_{s}:s\in S\}$, a history-dependent $S$-rectangular adversary policy $\kappa$ is a sequence of adversarial decision rules $\kappa=(\kappa_{0},\kappa_{1},\kappa_{2},\ldots)$. Each decision rule $\kappa_{t}$ specifies the conditional distribution of the next state given a history $h_{t}\in\mathbf{H}_{t}$ and an action $a\in A$, i.e., $\kappa_{t}(s^{\prime}|h_{t},a)\in[0,1]$ with $\sum_{s^{\prime}\in S}\kappa_{t}(s^{\prime}|h_{t},a)=1$. The history-dependent adversary policy class is
$\mathrm{K}_{\mathrm{H}}(\mathcal{P}):=\{\kappa=(\kappa_{0},\kappa_{1},\ldots):\kappa_{t}(\cdot|h_{t},\cdot)\in\mathcal{P}_{s_{t}},\ \text{where }h_{t}=(s_{0},a_{0},\ldots,s_{t}),\ \forall t\geq 0\}.$
@@ 3 Robust Bellman Equations and Optimality
In this section, we define the constant-gain robust Bellman equation and show that its solution determines the long-run average reward of the robust control problem. This also implies stationary optimality for the controller in the $\overline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}})$ and $\underline{\alpha}(\mu,\Pi_{\mathrm{H}},\mathrm{K}_{\mathrm{H}})$ case. The proofs of the results in this are deferred to Appendix A.
3.1 Robust Bellman Equation with a Constant Gain
@@
Similarly, $v^{\prime}_{\gamma}:S\rightarrow\mathbb{R}$ is the unique solution to the $\gamma$-discounted inf-sup equation if
$v_{\gamma}^{\prime}(s)=\inf_{p_{s}\in\mathcal{P}_{s}}\sup_{\phi\in\mathcal{Q}}E_{\phi,p_{s}}[r(s,A_{0})+\gamma v_{\gamma}^{\prime}(X_{1})],\quad\forall s\in S.$ (3.4)
@@
In the discounted setting, existence and uniqueness of solutions to (3.3) and (3.4) follow from a standard contraction mapping argument; see Wang et al., 2023a [24]. By contrast, in the average-reward setting, existing results establish the existence of a solution to (3.1) only in SA-rectangular settings and under additional assumptions [26, Theorem 8]. We substantially generalize these existence results to one–sided weak communication settings, which are robust analogues of the weakly communicating structures commonly assumed in classical MDPs to ensure constant-gain optimality [18].